### v1

@@ 4 Dense surface reconstruction solutions
SurfaceNet 2017 Point cloud Advantages: (1) High adaptability; (2) Efficient in processing. Disadvantages: (1) Dependent on training data; (2) Complexity of the learned features and their interaction with the MVS pipeline.
@@
NeuS 2021 Mesh Advantages: (1) Capable of achieving high-quality reconstruction; (2) Suitable for high-fidelity real-time rendering. Disadvantages: (1) Long training times; (2) High computational resource demands; (3) Sensitivity to data and parameter settings.
@@
SuGAR 2023 Mesh Advantages: (1) Improved computational efficiency; (2) Capable of providing high-quality detail representation. Disadvantages: (1) Immature in stability and robustness; (2) Sensitivity to data and parameter settings.
@@
4.1 Traditional multi-view stereo methods
@@
4.2 Deep learning-based multi-view stereo methods
@@
Hartmann et al. (Hartmann et al. (2017)) introduced a learnable multi-view cost metric to measure the multi-view photo-consistency between image patches. Subsequently, Ji et al. proposed SurfaceNet (Ji et al. (2017)), which learns cost volume regularization from geometry ground truth. Since then, various networks have been developed, with MVSNet (Yao et al. (2018)) being one of the most notable. MVSNet begins by extracting deep visual image features and then employs differentiable homography warping to build a 3D cost volume within the reference camera frustum. The method then applies regularization and refinement processes to produce the final output. MVSNet is also the first end-to-end method for learning depth map inference in MVS, and it has had a significant impact on subsequent approaches. Many later methods have built upon its framework. For example, Vis-MVSNet (Zhang et al. (2023)) extends MVSNet by formulating a more reliable cost volume to address the reconstruction of scenes with severe occlusions, achieving improved results. Some learning-based approaches have also sought to leverage the PatchMatch algorithm to avoid global cost volumes. Among these, PatchmatchNet (Wang et al. (2021a)) introduced the first end-to-end cascade formulation of PatchMatch. Later, inspired by the significant achievements of Vision Transformers (ViT) (Dosovitskiy (2020)) in various visual tasks, researchers have increasingly integrated transformers into MVS learning. MVSFormer (Cao et al. (2022)) represents a pioneering effort in this area, combining pre-trained ViTs for feature extraction with integrated architectures and training strategies, and has demonstrated promising results. Furthermore, methods such as TransMVSNet (Ding et al. (2022)) and MVSFormer++ (Cao et al. (2024)) have been developed. with MVSFormer++ achieving state-of-the-art results on the indoor DTU dataset (Aanæs et al. (2016)) and ranking top-1 on the outdoor Tanks-and-Temples dataset (Knapitsch et al. (2017)).
4.3 Neural radiance field-based methods
@@
4.4 Gaussian Splatting-based methods
@@
While Gaussians can achieve high-quality scene rendering, their application for surface extraction presents significant challenges and limitations that impact the accuracy and reliability of reconstruction. Issues include the unstructured nature of optimized Gaussians, multi-view inconsistencies, and conflicts between volumetric 3D Gaussians and the thin nature of surfaces (Guédon and Lepetit (2024); Huang et al. (2024a); Yu et al. (2024)). To address these challenges, researchers have proposed various solutions. SuGAR (Guédon and Lepetit (2024)) introduces a regularization term to better align Gaussians with the scene’s surface, and employs Poisson reconstruction to extract a mesh from the Gaussians, enabling accurate estimation of fine geometric details while being fast and user-friendly. NeuSG (Chen et al. (2023)) utilizes normal priors predicted by neural implicit models to refine the point cloud generated by 3D Gaussian Splatting for more accurate surface reconstruction. Huang et al. proposed 2D Gaussian Splatting (2D GS) (Huang et al. (2024a)), which uses 2D Gaussians for a more precise representation of the scene and employs TSDF fusion to reconstruct the mesh. Gaussian Opacity Fields (GOF)(Yu et al. (2024)) , derived from ray-tracing-based volume rendering of 3D Gaussians, directly extracts geometry from 3D Gaussians, eliminating the need for Poisson reconstruction (Kazhdan and Hoppe (2013a)) or TSDF fusion.
@@ 5.1 Experimental datasets
The real-world image data is provided by the Moorea Island Digital Ecosystem Avatar (IDEA) project. These images were captured underwater using a digital camera system that includes a PANASONIC LUMIX GH5S camera body, with a resolution of 3680×2760, and a wide-angle Lumix G 14mm f/2.5 lens. In this study, the data is organized into two datasets, named Coral-2018 and Coral-2019. The Coral-2018 dataset was collected in August 2018 and consists of 523 images, while the Coral-2019 dataset was collected in August 2019 and contains 318 images. All images in each dataset were taken along pre-planned paths, with an overlap rate between adjacent images ranging from approximately 70% to 85%, making them suitable for high-quality multi-view 3D reconstruction.
The synthetic datasets utilized in this study are created using the AirSim platform (Shah et al. (2018)), a simulator built on Unreal Engine (Sanders (2016)) that offers physically and visually realistic simulations. We imported models of coral reefs, rocks, and other elements into the simulation environments. Then, we applied the camera functionality to capture images of the coral reefs from various viewpoints. This approach yielded high-resolution, distortion-free images with pose ground truth, suitable for assessment purposes. Two synthetic datasets are designed for this study: Coral-UE4 and Coral-UE5, based on Unreal Engine 4 and Unreal Engine 5, respectively. The images in both datasets have a resolution of 2560×1440 pixels and were captured from a 360-degree perspective around the scenes. The Coral-UE4 dataset features simpler scenes with less texture, while the Coral-UE5 dataset includes more complex and richly textured environments. To evaluate the performance of different solutions in more challenging scenarios, both datasets are downsampled by uniformly selecting one-fourth of the images, resulting in lite datasets: Coral-UE4 (lite) with 17 images and Coral-UE5 (lite) with 18 images. The datasets and projects from this study are available at https://github.com/Atypical-Programmer/Coral3D.
@@ 5.2.1 Feature extraction and matching evaluation
Here we first evaluate their applications in camera pose estimation. For feature extraction, we consider two hand-crafted methods, SIFT (Lowe (2004)) and KAZE (Alcantarilla et al. (2012a)), as well as five deep learning-based approaches: SuperPoint (DeTone et al. (2018)), R2D2 (Revaud et al. (2019)), DISK (Tyszkiewicz et al. (2020)), ALIKED (Zhao et al. (2022)), and DeDoDe (Edstedt et al. (2024a)). It should be noted that DeDoDe V2 (Edstedt et al. (2024b)) has not yet been fully open sourced. The number of features extracted by each method is limited to 8000. For feature matching, two non-learning methods are employed: ratio-test (Lowe (2004)) and AdaLAM (Cavalli et al. (2020)). Additionally, three deep learning-based methods are used: SuperGlue (Sarlin et al. (2020)), LightGlue (Lindenberger et al. (2023)), and LoFTR (Sun et al. (2021)). Notably, LoFTR is a detector-free feature matching method that does not require keypoints. SuperGlue is available with two sets of pre-trained weights: one trained on the ScanNet dataset (Dai et al. (2017)) for indoor scenes, referred to as SG (indoor), and another trained on the MegaDepth dataset (Li and Snavely (2018)) for outdoor scenes, referred to as SG (outdoor). Since the data augmentation during the training process of SuperGlue incorporates SuperPoint, this method exhibits optimal performance when paired with SuperPoint. Similarly, pre-trained weights for LightGlue are available with SIFT, SuperPoint, DISK, and ALIKED features, all of which are included in the comparative analysis. LoFTR is implemented using the Python library “kornia” and provides three sets of pre-trained weights for correspondence search, i.e. “indoor model”, “new indoor model”, and “outdoor model”.
@@ 5.2.2 SfM reconstruction results evaluation
Based on the correspondences obtained from image matching, SfM reconstruction can be performed. This study utilizes COLMAP (Schonberger and Frahm (2016)) as the SfM framework, inputting the matching results from Section 5.2.1 into COLMAP for reconstruction. The number of features extracted from each image is still restricted to 8000. For LoFTR, the results from its outdoor model are input into DF-SfM He et al. (2024) for reconstruction. Additionally, the recently proposed end-to-end trained VGG-SfM Chen and Zhang (2019) is also evaluated. VGG-SfM supports the use of SIFT, SuperPoint, and ALIKED for extracting query points. However, due to its high memory consumption, VGG-SfM is not feasible for large-scale scenes like Coral-2018 and Coral-2019. In SfM reconstruction, a simple camera model with one focal length and two radial distortion parameters is applied to model the distortion.
@@
After completing the SfM reconstruction, we assess five metrics: Rate, Features, Points, Track, and $E_{rep}$. Rate represents the ratio of the number of images successfully registered to the total number of images. Features denotes the average number of successfully matched features on aligned images. Points refers to the number of 3D scene points generated by the SfM reconstruction (1k = 1000). Track indicates the average number of 2D features corresponding to a single 3D scene point, which can be understood as the average number of observations per point. $E_{rep}$ is the average reprojection error of matched features. Additionally, for the reconstruction results from the synthetic dataset, we align the reconstructed camera positions to the ground truth positions in the simulation environment using a seven-parameter transformation, and then calculate the position error $E_{loc}$ and the orientation error $E_{dir}$, where $E_{dir}$ denotes the angle between the reconstructed camera direction and the actual direction. The quantitative metrics for the real-world datasets are presented in Table 3. For the simulation dataset, Coral UE5 and Coral UE5 (lite) are taken as examples, as shown in Table 4.
@@
Table 3: Various metrics of the reconstruction results of Coral-2018 and Coral-2019 using different methods.
@@
Coral-2018 Coral-2019 Feature Match Rate Features Points Track ${E_{rep}}$ Rate Features Points Track ${E_{rep}}$ SIFT RT 498/523 3968.8 533,787 3.70 0.57 318/318 3946.6 324,741 3.86 0.51 AdaLAM 523/523 3391.0 364,790 4.86 0.75 318/318 3633.4 230,345 5.02 0.65 LightGlue 523/523 3887.0 367,532 5.53 1.07 318/318 4358.4 241,915 5.73 1.04 KAZE RT 467/523 3475.7 475,985 3.41 0.44 285/318 3324.0 302,276 3.13 0.33 AdaLAM 523/523 2713.4 308,286 4.60 0.64 314/318 2508.1 189,834 4.15 0.44 SuperPoint RT 458/523 3188.9 435,064 3.36 0.67 312/318 3226.1 277,544 3.63 0.64 AdaLAM 523/523 2952.5 387,553 3.98 0.84 316/318 3140.9 233,431 4.25 0.77 SG (indoor) 231/523 4093.4 244,768 3.86 0.83 300/318 3839.5 290,361 3.97 0.84 SG (outdoor) 523/523 4147.0 554,098 3.91 0.98 318/318 4569.2 354,975 4.09 0.94 LightGlue 523/523 2898.4 501,528 3.02 1.30 318/318 3006.2 306,996 3.11 1.35 R2D2 RT 100/523 1724.7 38,096 4.53 1.09 190/318 1090.9 51,964 3.99 1.01 AdaLAM 144/523 2228.6 65,801 4.88 1.23 283/318 2211.0 142,993 4.38 1.16 DISK RT 519/523 3497.5 438,245 4.14 1.04 307/318 3877.8 269,344 4.42 1.01 AdaLAM 520/523 3978.2 505,043 4.10 1.08 310/318 4330.9 306,455 4.38 1.06 LightGlue 46/523 5604.2 50,910 5.06 1.39 156/318 4556.0 177,453 4.01 1.32 ALIKED RT 523/523 2597.1 339,932 4.00 0.53 317/318 2804.3 213,382 4.17 0.45 AdaLAM 523/523 2955.5 393,632 3.93 0.60 317/318 3229.7 247,330 4.14 0.54 LightGlue 523/523 3427.2 488,691 3.67 1.22 318/318 4665.9 370,183 4.01 1.27 DeDoDe RT 519/523 2253.5 289,103 4.05 1.17 309/318 3120.2 221,701 4.35 1.23 AdaLAM 523/523 4198.3 478,774 4.59 1.33 317/318 4519.3 297,451 4.82 1.34 DF-SfM (LoFTR) 522/523 5065.5 1,158,799 2.28 0.45 312/318 5205.9 711,883 2.28 0.43
@@
Table 4: Various metrics of the reconstruction results of Coral-UE5 and Coral-UE5 (lite) using different methods.
@@
In general, R2D2 performs the worst, particularly on the Coral-2018 dataset, where only a small subset of the images are aligned. And it fails to achieve any meaningful reconstruction on the Coral-UE5 (lite) dataset. Additionally, in comparative analysis, R2D2 consistently shows the lowest values for Features and Points. This suboptimal performance is likely due to the R2D2 features’ significant degradation under moderate changes in viewpoint. In contrast, ALIKED exhibits significantly better performance. It achieves the highest overall Rate for reconstruction and demonstrates lower values for $E_{rep}$, $E_{loc}$, and $E_{dir}$, indicating that its results are not only complete and reliable but also accurate. SIFT generally produces favorable results, although its performance is suboptimal when combined with the ratio-test. In many cases, different approaches produce similar $E_{loc}$ and $E_{dir}$. This phenomenon can be attributed to the fact that most feature matching strategies can obtain sufficient correct correspondences under standard conditions, and the SfM pipeline’s outlier removal helps maintain reconstruction accuracy by eliminating significant outliers. However, the key challenge lies in maintaining this performance under challenging scenarios. When image overlap decreases, approaches that struggle to establish sufficient correct correspondences lead to either complete reconstruction failure or significantly increased errors, as evidenced by the performance degradation of multiple approaches on the Coral-UE5 (lite) dataset.
@@
In terms of feature matching, AdaLAM often proves beneficial in enhancing the completeness of the reconstruction. It effectively filters out significant errors and identifies potential correspondences, thereby increasing the Track value and contributing to a more stable SfM reconstruction. This is particularly relevant in real-world scenarios where the complexity of image acquisition often leads to challenging matching conditions and a higher likelihood of mismatches. Nevertheless, it is crucial to recognize that the use of AdaLAM may not necessarily improve accuracy. It can lead to an increase in reprojection error because the features matched by AdaLAM might not correspond precisely to the same 3D scene point but rather approximate locations. This can be considered as a trade-off between robustness and reliability versus precision.

### v2

@@ 2.3 Dense surface reconstruction
SurfaceNet 2017 Point cloud
Advantages: (1) High adaptability; (2) Efficient in processing.
Disadvantages: (1) Dependent on training data; (2) Complexity of the learned features and their interaction with the MVS pipeline.
@@
NeuS 2021 Mesh
Advantages: (1) Capable of achieving high-quality reconstruction; (2) Suitable for high-fidelity real-time rendering.
Disadvantages: (1) Long training times; (2) High computational resource demands; (3) Sensitivity to data and parameter settings.
@@
SuGAR 2023 Mesh
Advantages: (1) Improved computational efficiency; (2) Capable of providing high-quality detail representation.
Disadvantages: (1) Immature in stability and robustness; (2) Sensitivity to data and parameter settings.
@@
2.3.1 Traditional multi-view stereo methods
@@
2.3.2 Deep learning-based multi-view stereo methods
@@
Hartmann et al. (Hartmann et al. (2017)) introduced a learnable multi-view cost metric to measure the multi-view photo-consistency between image patches. Subsequently, Ji et al. proposed SurfaceNet (Ji et al. (2017)), which learns cost volume regularization from geometry ground truth. Since then, various networks have been developed, with MVSNet (Yao et al. (2018)) being one of the most notable. MVSNet begins by extracting deep visual image features and then employs differentiable homography warping to build a 3D cost volume within the reference camera frustum. The method then applies regularization and refinement processes to produce the final output. MVSNet is also the first end-to-end method for learning depth map inference in MVS, and it has had a significant impact on subsequent approaches. Many later methods have built upon its framework. For example, Vis-MVSNet (Zhang et al. (2023)) extends MVSNet by formulating a more reliable cost volume to address the reconstruction of scenes with severe occlusions, achieving improved results. Some learning-based approaches have also sought to leverage the PatchMatch algorithm to avoid global cost volumes. Among these, PatchmatchNet (Wang et al. (2021a)) introduced the first end-to-end cascade formulation of PatchMatch. Later, inspired by the significant achievements of Vision Transformers (ViT) (Dosovitskiy (2020)) in various visual tasks, researchers have increasingly integrated transformers into MVS learning. MVSFormer (Cao et al. (2022)) represents a pioneering effort in this area, combining pre-trained ViTs for feature extraction with integrated architectures and training strategies, and has demonstrated promising results. Furthermore, methods such as TransMVSNet (Ding et al. (2022)) and MVSFormer++ (Cao et al. (2024)) have been developed, with MVSFormer++ achieving state-of-the-art results on the indoor DTU dataset (Aanæs et al. (2016)) and ranking top-1 on the outdoor Tanks-and-Temples dataset (Knapitsch et al. (2017)).
2.3.3 Neural radiance field-based methods
@@
2.3.4 Gaussian Splatting-based methods
@@
While Gaussians can achieve high-quality scene rendering, their application for surface extraction presents significant challenges and limitations that impact the accuracy and reliability of reconstruction. Issues include the unstructured nature of optimized Gaussians, multi-view inconsistencies, and conflicts between volumetric 3D Gaussians and the thin nature of surfaces (Guédon and Lepetit (2024); Huang et al. (2024a); Yu et al. (2024)). To address these challenges, researchers have proposed various solutions. SuGAR (Guédon and Lepetit (2024)) introduces a regularization term to better align Gaussians with the scene’s surface, and employs Poisson reconstruction to extract a mesh from the Gaussians, enabling accurate estimation of fine geometric details while being fast and user-friendly. NeuSG (Chen et al. (2023)) utilizes normal priors predicted by neural implicit models to refine the point cloud generated by 3D Gaussian Splatting for more accurate surface reconstruction. Huang et al. proposed 2D Gaussian Splatting (2D GS) (Huang et al. (2024a)), which uses 2D Gaussians for a more precise representation of the scene and employs TSDF fusion to reconstruct the mesh. Gaussian Opacity Fields (GOF)(Yu et al. (2024)) , derived from ray-tracing-based volume rendering of 3D Gaussians, directly extracts geometry from 3D Gaussians, eliminating the need for Poisson reconstruction or TSDF fusion.
@@ 3.1 Experimental datasets
The real-world image data is provided by the Moorea Island Digital Ecosystem Avatar (IDEA) project. These images were captured underwater using a digital camera system that includes a PANASONIC LUMIX GH5S camera body, with a resolution of 3680×2760, and a wide-angle Lumix G 14mm f/2.5 lens. In this study, the data is organized into two datasets, named Coral-2018 and Coral-2019. The Coral-2018 dataset was collected in August 2018 and consists of 451 images, while the Coral-2019 dataset was collected in August 2019 and contains 318 images. All images in each dataset were taken along pre-planned paths, with an overlap rate between adjacent images ranging from approximately 70% to 85%, making them suitable for high-quality multi-view 3D reconstruction.
The synthetic datasets utilized in this study are created using the AirSim platform (Shah et al. (2018)), a simulator built on Unreal Engine (Sanders (2016)) that offers physically and visually realistic simulations. We imported models of coral reefs, rocks, and other elements into the simulation environments. Then, we applied the camera functionality to capture images of the coral reefs from various viewpoints. This approach yielded high-resolution, distortion-free images with pose ground truth, suitable for assessment purposes. Two synthetic datasets are created for this study: Coral-UE4 and Coral-UE5, based on Unreal Engine 4 and Unreal Engine 5, respectively. In Coral-UE4, the scenes are rendered without the inclusion of additional participating media, resulting in clean and unperturbed imaging conditions. In contrast, an underwater blueprint is used in Coral-UE5 to simulate the underwater environment. This system enables the inclusion of volumetric water effects such as light scattering, absorption, and attenuation, allowing for a more realistic approximation of the optical properties of water and their impact on image formation. The images in both datasets have a resolution of 2560×1440 pixels and were captured from a 360-degree perspective around the scenes. The Coral-UE4 dataset features simpler scenes with less texture, while the Coral-UE5 dataset includes more complex and richly textured environments. To evaluate the performance of different solutions in more challenging scenarios, both datasets are downsampled by uniformly sampling one-fourth of the images from the original dataset, resulting in lite datasets: Coral-UE4 (lite) with 17 images and Coral-UE5 (lite) with 18 images. The datasets and projects from this study are available at https://github.com/Atypical-Programmer/Coral3D.
@@ 3.2.1 Feature extraction and matching evaluation
Here we first evaluate their applications in camera pose estimation. For feature extraction, we consider two hand-crafted methods, SIFT (Lowe (2004)) and KAZE (Alcantarilla et al. (2012)), as well as five deep learning-based approaches: SuperPoint (DeTone et al. (2018)), R2D2 (Revaud et al. (2019)), DISK (Tyszkiewicz et al. (2020)), ALIKED (Zhao et al. (2022)), and DeDoDe (Edstedt et al. (2024a)). It should be noted that DeDoDe V2 (Edstedt et al. (2024b)) has not yet been fully open sourced. The number of features extracted by each method is limited to 8000. For feature matching, two non-learning methods are employed: ratio-test (Lowe (2004)) and AdaLAM (Cavalli et al. (2020)). Additionally, three deep learning-based methods are used: SuperGlue (Sarlin et al. (2020)), LightGlue (Lindenberger et al. (2023)), and LoFTR (Sun et al. (2021)). Notably, LoFTR is a detector-free feature matching method that does not require keypoints. SuperGlue is available with two sets of pre-trained weights: one trained on the ScanNet dataset (Dai et al. (2017)) for indoor scenes, referred to as SG (indoor), and another trained on the MegaDepth dataset (Li and Snavely (2018)) for outdoor scenes, referred to as SG (outdoor). Since the data augmentation during the training process of SuperGlue incorporates SuperPoint, this method exhibits optimal performance when paired with SuperPoint. Similarly, pre-trained weights for LightGlue are available with SIFT, SuperPoint, DISK, and ALIKED features, all of which are included in the comparative analysis. LoFTR is implemented using the Python library “kornia” and provides three sets of pre-trained weights for correspondence search, i.e. “indoor model”, “new indoor model”, and “outdoor model”.
@@ 3.2.2 SfM reconstruction results evaluation
Based on the correspondences obtained from image matching, SfM reconstruction can be performed. This study utilizes COLMAP (Schonberger and Frahm (2016)) as the SfM framework, inputting the matching results from Section 3.2.1 into COLMAP for reconstruction. The number of features extracted from each image is still restricted to 8000. For LoFTR, the results from its outdoor model are input into DF-SfM (He et al. (2024)) for reconstruction. Additionally, the recently proposed end-to-end trained VGG-SfM (Chen and Zhang (2019)) is also evaluated. VGG-SfM supports the use of SIFT, SuperPoint, and ALIKED for extracting query points. However, due to its high memory consumption, VGG-SfM is not feasible for large-scale scenes like Coral-2018 and Coral-2019. In SfM reconstruction, a simple camera model with one focal length and two radial distortion parameters is applied to model the distortion. DF-SfM and VGG-SfM are constrained by GPU memory and therefore resize the longer edge of each image to 1200 pixels prior to processing. For other methods, we use the original images without additional processing is applied, though different methods may apply their own image processing. In this study, we follow the default settings provided by each original implementation.
@@
Upon completion of SfM reconstruction, five evaluation metrics are computed: Rate, Features, Points, Track, and $E_{rep}$. Rate represents the ratio of the number of images successfully registered to the total number of images. Features denotes the average number of successfully matched features on aligned images. Points refers to the number of 3D scene points generated by the SfM reconstruction (1k = 1000). Track indicates the average number of 2D features corresponding to a single 3D scene point, which can be understood as the average number of observations per point. $E_{rep}$ is the average reprojection error of matched features.
For real-world datasets collected at Moorea Island (Coral-2018 and Coral-2019), five ground control points (GCPs) with planar photogrammetric coded targets were established, as illustrated in Figure 9. These targets, labelled Target 1 to Target 5, form the basis of an underwater geodetic control network. Their coordinates were determined using high-precision geodetic techniques (Nocerino et al. (2020)). Vertical heights were measured using geometric leveling with a green laser, while horizontal distances between points were carefully measured using a millimeter-graduated metal tape to support trilateration. The collected measurements were then processed through a network adjustment procedure to compute the final 3D coordinates of the GCPs. In this experiment, Target 5 is designated as a check point, and the remaining four as control points. Following initial SfM reconstruction, the GCP markers are identified across multiple images, enabling triangulation to estimate their positions in the arbitrary SfM coordinate system. To align the reconstructed model with the local geodetic control network, a seven-parameter transformation is estimated using the control points. Specifically, let $\mathbf{P}^{\mathrm{SfM}}_{\mathrm{ctrl},i}$ and $\mathbf{P}^{\mathrm{Local}}_{\mathrm{ctrl},i}$ denote the coordinates of the $i$-th control point in the SfM and local geodetic systems, respectively. The optimal transformation parameters, consisting of a scale factor $\hat{s}\in\mathbb{R}$, a rotation matrix $\hat{\mathbf{R}}\in\mathbb{R}^{3\times 3}$, and a translation vector $\hat{\mathbf{T}}\in\mathbb{R}^{3\times 1}$, are determined by solving the following least-squares minimization problem:
$(\hat{s},\hat{\mathbf{R}},\hat{\mathbf{T}})=\arg\min_{s,\mathbf{R},\mathbf{T}}\sum_{i\in C}\left\|s\,\mathbf{R}\,\mathbf{P}^{\mathrm{SfM}}_{\mathrm{ctrl},i}+\mathbf{T}-\mathbf{P}^{\mathrm{Local}}_{\mathrm{ctrl},i}\right\|^{2}_{2}$ (2)
where $C$ is the set of control point indices. The spatial accuracy in object space is then evaluated by computing RMSEs between the transformed SfM coordinates and the ground-truth positions. Here, $\left\|\cdot\right\|_{2}$ denotes the 3D Euclidean norm, i.e., the square root of the sum of squared coordinate differences. RMSEs are reported separately for the control points $E_{\mathrm{control}}$ and the check point $E_{\mathrm{check}}$, both in millimeters. For each point $j$, the point-wise error is defined as the Euclidean distance between its transformed SfM coordinate $\mathbf{P}^{\mathrm{SfM}}_{j}$ and its ground-truth coordinate $\mathbf{P}^{\mathrm{Local}}_{j}$:
$E_{j}=\left\|\hat{s}\,\hat{\mathbf{R}}\,\mathbf{P}^{\mathrm{SfM}}_{j}+\hat{\mathbf{T}}-\mathbf{P}^{\mathrm{Local}}_{j}\right\|_{2}$ (3)
For synthetic datasets, the reconstructed camera positions are aligned to the ground truth positions from the simulation environment using a seven-parameter transformation. From this alignment, we compute the root mean square errors (RMSE) for camera positions $E_{loc}$ and orientations $E_{dir}$. The orientation errors are calculated based on the angular deviations between the reconstructed and ground-truth camera directions, measured in degrees. The quantitative metrics for the real-world datasets are presented in Table 3. For the simulation datasets, Coral UE5 and Coral UE5 (lite) are taken as examples, as shown in Table 4.
Figure 9: The distribution of GCPs in the Moorea Island survey area.
@@
Table 3: Various metrics of the reconstruction results of Coral-2018 and Coral-2019 using different methods. The missing values, indicated by dashes, are due either to reconstruction failure or to the inapplicability of certain metrics to the corresponding methods.
@@
Coral-2018 Coral-2019 Feature Match Rate Features Points Track $E_{rep}$} $E_{control}$} $E_{check}$} Rate Features Points Track $E_{rep}$} $E_{control}$} $E_{check}$} SIFT RT 451/451 4244.8 503,475 3.80 0.56 7.99 9.74 318/318 3946.6 324,741 3.86 0.51 3.35 41.37 AdaLAM 451/451 3799.5 338,350 5.06 0.72 8.05 10.71 318/318 3633.4 230,345 5.02 0.65 3.42 41.65 LightGlue 451/451 4457.1 338,517 5.94 1.07 7.73 11.84 318/318 4358.4 241,915 5.73 1.04 3.51 43.16 KAZE RT 447/451 3534.1 454,122 3.48 0.38 8.94 8.71 285/318 3324.0 302,276 3.13 0.33 - - AdaLAM 451/451 2924.1 272,200 4.84 0.51 9.08 10.54 314/318 2508.1 189,834 4.15 0.44 4.16 36.15 SuperPoint RT 404/451 3426.8 400,455 3.46 0.65 57.39 70.00 312/318 3226.1 277,544 3.63 0.64 4.13 34.23 AdaLAM 451/451 3173.0 344,664 4.15 0.80 25.58 25.64 316/318 3140.9 233,431 4.25 0.77 4.40 36.81 SG (indoor) 229/451 4127.5 244,283 3.87 0.83 - - 300/318 3839.5 290,361 3.97 0.84 - - SG (outdoor) 451/451 4777.7 527,298 4.09 0.97 10.72 5.54 318/318 4569.2 354,975 4.09 0.94 3.75 36.49 LightGlue 451/451 3654.5 529,035 3.12 1.35 8.48 5.41 318/318 3006.2 306,996 3.11 1.35 4.98 40.90 R2D2 RT 100/451 1724.1 38,071 4.53 1.09 - - 190/318 1090.9 51,964 3.99 1.01 19.30 31.98 AdaLAM 122/451 2530.8 61,246 5.04 1.23 16.81 16.90 283/318 2211.0 142,993 4.38 1.16 16.92 45.54 DISK RT 448/451 3826.8 394,511 4.35 1.02 7.39 3.81 307/318 3877.8 269,344 4.42 1.01 7.05 33.23 AdaLAM 448/451 4369.5 454,828 4.30 1.07 8.03 2.26 310/318 4330.9 306,455 4.38 1.06 5.04 34.29 LightGlue 46/451 5605.0 50,932 5.06 1.39 - - 156/318 4556.0 177,453 4.01 1.32 - - ALIKED RT 451/451 2968.4 319,944 4.18 0.50 8.45 4.13 317/318 2804.3 213,382 4.17 0.45 3.94 40.57 AdaLAM 451/451 3363.8 365,819 4.15 0.58 9.99 4.12 317/318 3229.7 247,330 4.14 0.54 3.63 40.59 LightGlue 451/451 4231.7 498,845 3.83 1.25 11.02 6.13 318/318 4665.9 370,183 4.01 1.27 4.23 37.91 DeDoDe RT 447/451 2465.0 260,041 4.24 1.18 30.97 54.90 309/318 3120.2 221,701 4.35 1.23 5.07 38.64 AdaLAM 451/451 4654.1 426,499 4.92 1.36 42.32 22.33 317/318 4519.3 297,451 4.82 1.34 4.15 41.65 DF-SfM (LoFTR) 451/451 5193.0 1,019,030 2.30 0.45 37.92 52.55 312/318 5205.9 711,883 2.28 0.43 7.65 30.17
@@
Table 4: Various metrics of the reconstruction results of Coral-UE5 and Coral-UE5 (lite) using different methods. The missing values, indicated by dashes, are due either to reconstruction failure or to the inapplicability of certain metrics to the corresponding methods.
@@
In general, R2D2 exhibits the weakest performance, particularly on the real-world datasets, where many images fail to be successfully aligned and the resulting errors remain at the centimeter level, falling short of the sub-centimetric accuracy for precise modeling. It also fails to achieve any meaningful reconstruction on the Coral-UE5 (lite) dataset. In comparative analyses, R2D2 consistently yields the lowest values for Features and Points. This suboptimal performance is likely due to the R2D2 features’ significant degradation under moderate changes in viewpoint. In contrast, ALIKED exhibits significantly better performance. It achieves the highest overall Rate for reconstruction and demonstrates lower values across various error metrics, indicating that its results are not only complete and reliable but also accurate. SIFT generally produces favorable results, although its performance is suboptimal when handling Coral-UE5 (lite) with the ratio-test. For synthetic datasets, different approaches often produce similar $E_{loc}$ and $E_{dir}$ because most feature matching strategies can obtain sufficient correct correspondences under standard conditions. Furthermore, the SfM pipeline’s outlier removal helps maintain reconstruction accuracy by eliminating significant outliers, and performance differences become apparent only when image overlap decreases. Approaches struggling to establish an adequate number of correct correspondences lead to either complete reconstruction failure or significantly increased errors, as evidenced by the performance degradation of multiple approaches on the Coral-UE5 (lite) dataset. For more complex real-world datasets, SIFT, DISK and ALIKED achieve millimeter-level accuracy on Coral-2018, as reflected in their low $E_{control}$ and $E_{check}$. SuperPoint achieves high accuracy only when paired with learning-based matching methods and otherwise suffers from large errors. DeDoDe produces large errors despite performing well on synthetic datasets, suggesting limited robustness for practical high-precision coral reef modeling. As for Coral-2019, the sparse coverage of the check point in the collected imagery results in generally larger errors. Nevertheless, the residuals at the control points reveal that most methods perform comparably, except for R2D2, whose errors are the largest and render it nearly unusable.
@@
In terms of feature matching, AdaLAM often proves beneficial in enhancing the completeness of the reconstruction. It effectively filters out significant errors and identifies potential correspondences, thereby increasing the Track value and contributing to a more stable SfM reconstruction. This is particularly relevant in real-world scenarios where the complexity of image acquisition often leads to challenging matching conditions and a higher likelihood of mismatches. Nevertheless, it is crucial to recognize that the use of AdaLAM does not necessarily improve accuracy and may, in fact, slightly increase the error in many cases. It can lead to an increase in reprojection error because the features matched by AdaLAM might not correspond precisely to the same 3D scene point but rather approximate locations. This can be considered as a trade-off between robustness and reliability versus precision.